### v4

# Sum of Two Squares in Biquadratic Fields

Abstract
This paper gives an algorithm to determine whether a number in a biquadratic field is a sum of two squares, based on local-global principle of isotropy of quadratic forms.
Keywords: Biquadratic field, Quadratic Forms, Isotropy
Introduction
Let $F_{1}=\mathbb{Q}(\sqrt{a})$, $F_{2}=\mathbb{Q}(\sqrt{b})$ be two quadratic fields, with $a$, $b$ square-free integers and $ab$ non-square. Put $c=\frac{ab}{gcd(a,b)^{2}}$ and $F=F_{1}F_{2}$, $F_{3}=\mathbb{Q}(\sqrt{c})$, then $c$ is also square-free, and $F_{1}$, $F_{2}$ and $F_{3}$ are the three quadratic subfields of $F$. Let $\sigma_{i}$ be the elements of $Gal(F/\mathbb{Q})\simeq(\mathbb{Z}/2)^{2}$ other than $id$ satisfying $\sigma_{i}|_{F_{i}}=id|_{F_{i}}$.
Some critical values about $F$, including integers in $F$ and discriminant of $F$, are already given in [1].
In this article, using local-global principle of isotropy of quadratic forms (see 66:1 of [2]), we will give an explicit algorithm to determine whether a number $s_{0}+s_{1}\sqrt{a}+s_{2}\sqrt{b}+s_{3}\sqrt{c}$ ($s_{0},s_{1},s_{2},s_{3}\in\mathbb{Z}$) is a sum of two squares in ${F}$. First we prepare some lemmas for convenience (see [4]):
Lemma 0.1.
Let $\mathbb{Q}_{2}(\sqrt{k})$ be a quadratic extension on $\mathbb{Q}_{2}$, $k\in\{2,3,7,10,11,14,5\}$, then $t\in\mathbb{Q}_{2}(\sqrt{k})$ is a square, if and only if
($k=2$)$t=2^{m}r$, $m\geq 0$ an integer, $\sqrt{2}\nmid r$, $r\equiv 1,3+2\sqrt{2}\pmod{4\sqrt{2}}$.
($k=3$)$t=(\sqrt{3}-1)^{2m}r$, $m\geq 0$ an integer, $(\sqrt{3}-1)\nmid r$, $r\equiv 1,3\pmod{4(\sqrt{3}-1)}$.
($k=7$)$t=(3-\sqrt{7})^{2m}r$, $m\geq 0$ an integer, $(3-\sqrt{7})\nmid r$, $r\equiv\pm 1\pmod{4(3-\sqrt{7})}$.
($k=11$)$t=(\sqrt{11}-3)^{2m}r$, $m\geq 0$ an integer, $(\sqrt{11}-3)\nmid r$, $r\equiv 1,3\pmod{4(\sqrt{11}-3)}$.
($k=14$)$t=(4-\sqrt{14})^{2m}r$, $m\geq 0$ an integer, $(4-\sqrt{14})\nmid r$, $r\equiv\pm 1\pmod{4(4-\sqrt{14})}$.
($k=10$)$t=Mr$, $v_{(2,\sqrt{10})}(M)$ is even, $v_{(2,\sqrt{10})}(r)=0$, $r\equiv 1,3+2\sqrt{10}\pmod{4(2,\sqrt{10})}$.
($k=5$)$t=4^{m}r$, $m\geq 0$ an integer, $2\nmid r$, $r\equiv 1,\frac{3\pm\sqrt{5}}{2}\pmod{4}$.
Lemma 0.2.
Take assumptions as the last lemma and let $h$ be a unit in $O_{\mathbb{Q}_{2}(\sqrt{k})}$, then $h$ is the sum of two squares if and only if
[…]
($k=5$)$h\equiv 1,3,\frac{\pm 3\pm\sqrt{5}}{2}\pmod{4}$.
1 Prime Numbers
First we state some facts. For any odd prime number $p$, either $p\nmid abc$, or $p$ exactly divides two of $\{a,b,c\}$. In the former case $(\frac{a}{p})(\frac{b}{p})=(\frac{c}{p})$ ($\frac{*}{*}$ denotes Legendre symbol), and in the latter case, say $p|a$ and $p|b$, $(\frac{a/p}{p})(\frac{b/p}{p})=(\frac{c}{p})$.
It is well-known that a prime number $q$ is a sum of two squares in $\mathbb{Q}$ if and only if it $q\not\equiv 3\pmod{4}$. Let $q$ be a prime number with $q\equiv 3\pmod{4}$. Next we find all biquadratic fields where $q$ becomes a sum of two squares.
Lemma 1.1.
(1)-1 is a sum of two squares in $F$ if $F$ is imaginary, unless $a$ and $b$ are both $\equiv 1\pmod{8}$.
(2)$q$ is a sum of two squares in $F$ unless $(\frac{a}{q})=(\frac{b}{q})=1$ or $a,b\equiv 1\pmod{8}$.
Proof.
(1)We only need $(\frac{-1,-1}{\mathcal{P}})=1$ for $\mathcal{P}$ infinite or dyadic. By lemma 0.2, $(\frac{-1,-1}{\mathcal{P}})=1$ if and only if $F_{\mathcal{P}}\not\simeq\mathbb{Q}_{2}$, i.e, $2O_{F}$ does not completely split.
(2)Of course $(\frac{-1,q}{\mathcal{P}})=1$ at infinite places. We only need to discuss $q$-adic and dyadic cases. By lemma 0.2, $(\frac{-1,q}{\mathcal{P}})=1$($\mathcal{P}$ dyadic) if and only if $2O_{F}$ does not completely split. In this case, if $(\frac{-1,q}{\mathcal{Q}})=-1$ for $q$-adic place $\mathcal{Q}$, then $F_{\mathcal{Q}}$ must not contain $\mathbb{Q}_{q}(\sqrt{-1})$, $\mathbb{Q}_{q}(\sqrt{q})$ or $\mathbb{Q}_{q}(\sqrt{-q})$ (in the last case $(\frac{-1,-1}{\mathcal{Q}})=1$ , making $(\frac{-1,q}{\mathcal{Q}})=(\frac{-1,-1}{\mathcal{Q}})(\frac{-1,-q}{\mathcal{Q}})=1$). So $F_{\mathcal{Q}}\simeq\mathbb{Q}_{q}$, i.e, $(\frac{a}{q})=(\frac{b}{q})=1$. ∎
2 General cases
Let $S=s_{0}+s_{1}\sqrt{a}+s_{2}\sqrt{b}+s_{3}\sqrt{c}\in F$ in introduction. By 66:1 and 63:12 of [2], to decide whether $S$ is a sum of squares in $F$, we only need to calculate $(\frac{-1,S}{\mathcal{P}})$s for all infinite, dyadic or $p$-adic places for $p\equiv 3\pmod{4}$ and $p|N_{F/\mathbb{Q}}(S)$. The infinite case is obvious: $(\frac{-1,S}{\mathcal{\infty}})=-1$ if and only if $F$ is real and $S$ is not totally positive.
Let $\mathcal{P}\cap\mathbb{Z}=p\equiv 3\mod 4$, $p|N_{F/\mathbb{Q}}(S)$.
2.1 Algorithm about unramified $p$
First we talk about the case $p\nmid abc$. If $(\frac{a}{p})=(\frac{b}{p})=(\frac{c}{p})=1$, then $pO_{F}$ completely splits, and four $(F_{\mathcal{P}})$s are isomorphic to $\mathbb{Q}_{p}$. set
$T_{i}=S\sigma_{i}(S)(i=1,2,3)$
Then $T_{i}\in F_{i}$, and $S$ is a sum of two squares implies that so are all $T_{i}$s in $F_{i}$s. Conversely, if all $T_{i}$s are sum of two squares, then $(\frac{-1,T_{i}}{\mathcal{P}})$s ($\mathcal{P}$ finite and non-dyadic) are all 1, which implies that for every $p$-adic places with $p\equiv 3\pmod{4}$, $F_{\mathcal{P}}$ contains a subfield isomorphic to $\mathbb{Q}_{p}(\sqrt{-1})$, or $v_{\mathcal{P}}(T_{i})$s are all even.
We need to state some facts about $p$-adic ideals. Let $A,B,C$ be (mod $p$) square roots of $a,b,c$ respectively, satisfying $C\mathrm{gcd}(a,b)\equiv AB\pmod{p}$, then $\sqrt{c}-C\in(p,\sqrt{a}-A,\sqrt{b}-B)$. Take
$\mathcal{P}_{1}=(p,\sqrt{a}-A,\sqrt{b}-B,\sqrt{c}-C)$
$\mathcal{P}_{2}=(p,\sqrt{a}-A,\sqrt{b}+B,\sqrt{c}+C)$
$\mathcal{P}_{3}=(p,\sqrt{a}+A,\sqrt{b}-B,\sqrt{c}+C)$
$\mathcal{P}_{4}=(p,\sqrt{a}+A,\sqrt{b}+B,\sqrt{c}-C)$
Then
$pO_{F_{1}}=\mathfrak{p}_{11}\mathfrak{p}_{12},\mathfrak{p}_{11}O_{F}=(p,\sqrt{a}-A)O_{F}=\mathcal{P}_{1}\mathcal{P}_{2},\mathfrak{p}_{12}O_{F}=(p,\sqrt{a}+A)O_{F}=\mathcal{P}_{3}\mathcal{P}_{4}$
$pO_{F_{2}}=\mathfrak{p}_{21}\mathfrak{p}_{22},\mathfrak{p}_{21}O_{F}=(p,\sqrt{b}-B)O_{F}=\mathcal{P}_{1}\mathcal{P}_{3},\mathfrak{p}_{22}O_{F}=(p,\sqrt{b}+B)O_{F}=\mathcal{P}_{2}\mathcal{P}_{4}$
$pO_{F_{1}}=\mathfrak{p}_{11}\mathfrak{p}_{12},\mathfrak{p}_{31}O_{F}=(p,\sqrt{c}-C)O_{F}=\mathcal{P}_{1}\mathcal{P}_{4},\mathfrak{p}_{32}O_{F}=(p,\sqrt{c}+C)O_{F}=\mathcal{P}_{2}\mathcal{P}_{3}$
Thus we have
$2v_{\mathcal{P}_{1}}(S)$ $=$ $v_{\mathcal{P}_{1}}(S\sigma_{1}(S))+v_{\mathcal{P}_{1}}(S\sigma_{2}(S))-v_{\mathcal{P}_{1}}(\sigma_{1}(S)\sigma_{2}(S))$
$=$ $v_{\mathfrak{p}_{11}}(S\sigma_{1}(S))+v_{\mathfrak{p}_{21}}(S\sigma_{2}(S))-v_{\mathfrak{p}_{31}}(\sigma_{1}(S)\sigma_{2}(S))$
Similarly, we have
$2v_{\mathcal{P}_{2}}(S)=v_{\mathfrak{p}_{11}}(S\sigma_{1}(S))+v_{\mathfrak{p}_{22}}(S\sigma_{2}(S))-v_{\mathfrak{p}_{32}}(\sigma_{1}(S)\sigma_{2}(S))$
$2v_{\mathcal{P}_{3}}(S)=v_{\mathfrak{p}_{12}}(S\sigma_{1}(S))+v_{\mathfrak{p}_{21}}(S\sigma_{2}(S))-v_{\mathfrak{p}_{32}}(\sigma_{1}(S)\sigma_{2}(S))$
$2v_{\mathcal{P}_{4}}(S)=v_{\mathfrak{p}_{12}}(S\sigma_{1}(S))+v_{\mathfrak{p}_{22}}(S\sigma_{2}(S))-v_{\mathfrak{p}_{31}}(\sigma_{1}(S)\sigma_{2}(S))$
Therefore given an algorithm to calculate powers of a number in a prime ideal of a quadratic field, that in a biquadratic field can be obtained.
Remark.
Here is an algorithm to calculate the power of a number in an explicit ideal of a quadratic field:
Take $\mathfrak{p}_{11}=(p,\sqrt{a}-A)$ as above, with $pO_{F_{1}}$ splits and $r=x+y\sqrt{a}$, $x,y\in\mathbb{Z}$. If $p\nmid\mathrm{gcd}(x,y)$, then replace $\sqrt{a}$ to $A$ in $r$ we get $\hat{r}=x+yA$. If $p|\hat{r}$ then $v_{\mathfrak{p}_{11}}(r)=v_{p}(x^{2}-ay^{2})$, otherwise $v_{\mathfrak{p}_{11}}(r)=0$. If $p^{f}||\mathrm{gcd}(x,y)$, then $v_{\mathfrak{p}_{11}}(r)=f+v_{\mathfrak{p}_{11}}(\frac{r}{p^{f}})$.
Lemma 2.1.
If $p\equiv 3\pmod{4}$ is a prime dividing $N_{F/\mathbb{Q}}(S)$ with $(\frac{a}{p})=(\frac{b}{p})=(\frac{c}{p})=1$, then $(\frac{-1,S}{\mathcal{P}})=1$ for all four $p$-adic places $\mathcal{P}$, if and only if
$v_{\mathfrak{p}_{11}}(S\sigma_{1}(S))+v_{\mathfrak{p}_{21}}(S\sigma_{2}(S))-v_{\mathfrak{p}_{31}}(\sigma_{1}(S)\sigma_{2}(S))$
$v_{\mathfrak{p}_{11}}(S\sigma_{1}(S))+v_{\mathfrak{p}_{22}}(S\sigma_{2}(S))-v_{\mathfrak{p}_{32}}(\sigma_{1}(S)\sigma_{2}(S))$
$v_{\mathfrak{p}_{12}}(S\sigma_{1}(S))+v_{\mathfrak{p}_{21}}(S\sigma_{2}(S))-v_{\mathfrak{p}_{32}}(\sigma_{1}(S)\sigma_{2}(S))$
$v_{\mathfrak{p}_{12}}(S\sigma_{1}(S))+v_{\mathfrak{p}_{22}}(S\sigma_{2}(S))-v_{\mathfrak{p}_{31}}(\sigma_{1}(S)\sigma_{2}(S))$
are all divided by 4.
If $(\frac{a}{p}),(\frac{b}{p}),(\frac{c}{p})$ is an arrangement of -1,-1,1 with $p\equiv 3\pmod{4}$ then $F_{\mathcal{P}}$s are all isomorphic to $\mathbb{Q}_{p}(\sqrt{-1})$, hence all $(\frac{-1,S}{\mathcal{P}})=1$.
2.2 Algorithm about ramified $p$
First set $p|abc$, say, $p|a$ and $p|b$.
If $(\frac{a/p}{p})\neq(\frac{b/p}{p})$, then $(\frac{c}{p})=-1$, so $F_{\mathcal{P}}$ has a subfield isomorphic to $\mathbb{Q}_{p}(\sqrt{-1})$, making -1 a square and thus $(\frac{-1,S}{\mathcal{P}})=1$.
Assume $(\frac{a/p}{p})=(\frac{b/p}{p})$. Then we have explicit expression of $p$-adic ideals:
$pO_{F_{1}}=\mathfrak{p}_{1}^{2},\mathfrak{p}_{1}O_{F}=\mathcal{P}_{1}\mathcal{P}_{2}$
$pO_{F_{2}}=\mathfrak{p}_{2}^{2},\mathfrak{p}_{2}O_{F}=\mathcal{P}_{1}\mathcal{P}_{2}$
$pO_{F_{3}}=\mathfrak{p}_{31}\mathfrak{p}_{32},\mathfrak{p}_{3i}O_{F}=\mathcal{P}_{i}^{2}$
with $i\in\{1,2\}$, and
$\mathfrak{p}_{1}=(p,\sqrt{a}),\mathfrak{p}_{2}=(p,\sqrt{b}),\mathfrak{p}_{3i}=(p,\sqrt{c}+(-1)^{i}C)$
where $C^{2}\equiv c\pmod{p}$, and
$\mathcal{P}_{i}=(p,\sqrt{a},\sqrt{b},\sqrt{c}+(-1)^{i}C).$
Thus in order to determine $v_{\mathcal{P}_{i}}(S)$, we denote
$2v_{\mathcal{P}_{1}}(S)$ $=$ $v_{\mathcal{P}_{1}}(S\sigma_{1}(S))+v_{\mathcal{P}_{1}}(S\sigma_{2}(S))-v_{\mathcal{P}_{1}}(\sigma_{1}(S)\sigma_{2}(S))$
$=$ $v_{\mathfrak{p}_{1}}(S\sigma_{1}(S))+v_{\mathfrak{p}_{2}}(S\sigma_{2}(S))-2v_{\mathfrak{p}_{31}}(\sigma_{1}(S)\sigma_{2}(S))$
Similarly we have
$2v_{\mathcal{P}_{2}}(S)=v_{\mathfrak{p}_{1}}(S\sigma_{1}(S))+v_{\mathfrak{p}_{2}}(S\sigma_{2}(S))-2v_{\mathfrak{p}_{32}}(\sigma_{1}(S)\sigma_{2}(S)).$
Remark.
One can easily prove that if $x,y\in\mathbb{Z}$, $p^{f}||\mathrm{gcd}(x,y)$, then
$v_{\mathfrak{p}_{1}}(x+y\sqrt{a})=f+v_{p}(\frac{x^{2}-ay^{2}}{p^{2f}})=v_{p}(x^{2}-ay^{2})-f.$
Thus we obtain that
Lemma 2.2.
If $p\equiv 3\pmod{4}$ is a prime dividing $N_{F/\mathbb{Q}}(S)$, and $abc$ (say, $p|a$ and $p|b$), and $(\frac{a/p}{p})=(\frac{b/p}{p})$ (equivalent that $(\frac{c}{p})=1$), then $(\frac{-1,S}{\mathcal{P}})=1$ for both $p$-adic places $\mathcal{P}$, if and only if both $v_{\mathfrak{p}_{1}}(S\sigma_{1}(S))+v_{\mathfrak{p}_{2}}(S\sigma_{2}(S))-2v_{\mathfrak{p}_{3i}}(\sigma_{1}(S)\sigma_{2}(S))$ are divided by 4 for $i\in\{1,2\}$.
2.3 Dyadic cases
Finally we need to compute $(\frac{-1,S}{\mathcal{P}})$ for dyadic places $\mathcal{P}$. (Of course we assume $(\frac{-1,S}{\mathcal{P}})=1$ all infinite and non-dyadic places $\mathcal{P}$.) If none of $a,b,c$ is $\equiv 1\pmod{8}$ (For example, $(a,b,c)=(2,3,6)$ or $(2,5,10)$), then $F$ has only one dyadic place, and by Hilbert Reciprocity Law we have $(\frac{-1,S}{\mathcal{P}})=1$.
Next we assume one of $\{a,b,c\}$, say $a$, is $\equiv 1\pmod{8}$. Then $b\equiv c\pmod{8}$.
(1)If $b\not\equiv 1\pmod{4}$, i.e, $2O_{F}$ splits and ramifies, then $F_{\mathcal{P}}$s are isomorphic to a quadratic ramified extension of $\mathbb{Q}_{2}$. By lemma 0.2, $(\frac{-1,S}{\mathcal{P}})=1$ for both dyadic places $\mathcal{P}$, if and only if $S=2^{f}(2S^{\prime}+1)$, with $f=v_{2}(S)$ and $S^{\prime}\in O_{F}$. (Recall that algebraic integers in biquadratic fields are given in Theorem 2 of [1], e.g, $\frac{1}{4}(1-\sqrt{21}+\sqrt{33}+\sqrt{77})$ is an algebraic integer in $\mathbb{Q}(\sqrt{21},\sqrt{33})$.)
(2)If $b\equiv 5\pmod{8}$, i.e, $2O_{F}$ splits and inerts, then both $F_{\mathcal{P}}$s are isomorphic to $\mathbb{Q}_{2}(\sqrt{5})$. Let $S^{\prime\prime}=\frac{S}{2^{v_{2}(S)}}$($0\leq v_{2}(S)\leq 2$) and
$e(N)=\left\{\begin{array}[]{lll}E&&if\ N\equiv E^{2}\pmod{256}\ odd\ and\ square-free,1\leq E(W)\leq 63\\ \sqrt{5}&&if\ N\equiv 5\pmod{32}\\ \sqrt{5}(1+2^{2}+2^{3})&&if\ N\equiv 13\pmod{32}\\ \sqrt{5}(1+2^{3})&&if\ N\equiv 21\pmod{32}\\ \sqrt{5}(1+2^{2})&&if\ N\equiv 29\pmod{32}\\ s_{1}s_{2}\dots s_{g}e(W)&&if\ w=s_{1}^{2}s_{2}^{2}\dots s_{g}^{2}W,W\equiv 1\pmod{4}\ square-free\end{array}\right.$
(Recall Lemma 2.7 of [4].) Replace $\sqrt{b}$ with $e(b)$, $\sqrt{c}$ with $e(c)$, and $\sqrt{a}$ with $A$ that $A\mathrm{gcd}(b,c)\equiv e(b)e(c)\pmod{16}$, then in $F_{\mathcal{P}}\simeq\mathbb{Q}_{2}(\sqrt{5})$ we have
$S^{\prime\prime}\equiv\frac{1}{2^{v_{2}(S)}}(s_{0}+s_{1}A+s_{2}e(b)+s_{3}e(c))\pmod{4}$
Finally, $\frac{1}{2^{v_{2}(S)}}(s_{0}+s_{1}A+s_{2}e(b)+s_{3}e(c))$ and then $S^{\prime\prime}$ can be decided a sum of two squares or not in $\mathbb{Q}_{2}(\sqrt{5})$ by Lemma 0.2.
(3)If $b\equiv 1\pmod{8}$, i.e, $2O_{F}$ completely splits, then $0\leq v_{2}(S)\leq 2$. Take $S^{\prime\prime}=\frac{S}{v_{2}(S)}$, where replacing $\sqrt{b}$ with $e(b)$, $\sqrt{c}$ with $e(c)$, and $\sqrt{a}$ with $A$ that $A\mathrm{gcd}(b,c)\equiv e(b)e(c)\pmod{16}$, then in $F_{\mathcal{P}}\simeq\mathbb{Q}_{2}$ we have
$S^{\prime\prime}\equiv\frac{1}{2^{v_{2}(S)}}(s_{0}+s_{1}A+s_{2}e(b)+s_{3}e(c))\pmod{4}.$
Finally, $\frac{1}{2^{v_{2}(S)}}(s_{0}+s_{1}A+s_{2}e(b)+s_{3}e(c))$ and then $S^{\prime\prime}$ can be decided a sum of two squares or not in $\mathbb{Q}_{2}$ by this well-known proposition: A unit in $O_{\mathbb{Q}_{2}}$ is a sum of two squares if and only if its odd part is $\equiv 1\pmod{4}$.
Concluding all process above, and 66:1 of [1], we conclude that
Theorem 2.3.
Let $S=s_{0}+s_{1}\sqrt{a}+s_{2}\sqrt{b}+s_{3}\sqrt{c}$ be a number in a biquadratic field $F=\mathbb{Q}(\sqrt{a},\sqrt{b})$, where $a,b$ are square-free integers and $ab$ is not a square, $c=\frac{ab}{\mathrm{gcd}(a,b)^{2}}$, and $(s_{0},s_{1},s_{2},s_{3})=1$, then $S$ is a sum of two squares in $F$ if and only if
(1)$S$ is totally positive if $F$ is a real field;
(2)For every prime number $p|N_{F/\mathbb{Q}}(S)$ with $p\equiv 3\pmod{4}$ and $(\frac{a}{p})=(\frac{b}{p})=1$, conditions in Lemma 2.1 are satisfied.
(3)For every prime number $p|N_{F/\mathbb{Q}}(S)$ with $p\equiv 3\pmod{4}$ and $p|abc$, say $p|a$ and $p|b$, and $(\frac{c}{p})=1$, conditions in Lemma 2.2 are satisfied.
(4)If at least one of $a,b,c$ is $\equiv 1\pmod{8}$, say, $a\equiv 1(mod8)$,
(4-1)For $b\not\equiv 1\pmod{4}$, there is an $S^{\prime}\in O_{F}$ satisfying $S=2^{f}(2S^{\prime}+1)$, $f\in\mathbb{Z}$;
(4-2)For $b\equiv 5\pmod{8}$, $s_{0}+s_{1}A+s_{2}e(b)+s_{3}e(c)=2^{f}S^{\prime\prime}$, where $S^{\prime\prime}\equiv 1,3,\frac{\pm 3\pm\sqrt{5}}{2}\pmod{4}$, $f=v_{2}(S)$;
(4-3)For $b\equiv 1\pmod{8}$, $s_{0}+s_{1}A+s_{2}e(b)+s_{3}e(c)=2^{f}S^{\prime\prime}$, where $S^{\prime\prime}\equiv 1\pmod{4}$, $f=v_{2}(S)$.
Finally, for most general cases that $(s_{0},s_{1},s_{2},s_{3})=1$, we have that
Theorem 2.4.
Take assumptions in Theorem 2.3. Let $\mathcal{S}=Q^{2}UVS$, $0\neq Q\in\mathbb{Q}$, $U$(resp. $V$) is the product of distinct prime numbers that $\equiv 1,2$(resp. $3$)$\pmod{4}$, then $\mathcal{S}$ is a sum of two squares in $F$ if and only if
(1)$\mathcal{S}$ is totally positive if $F$ is a real field;
(2)For every prime number $p|N_{F/\mathbb{Q}}(S)$ with $p\equiv 3\pmod{4}$ and $(\frac{a}{p})=(\frac{b}{p})=1$, either the four values in Lemma 2.1 are all $\equiv 0\pmod{4}$ with $p\nmid V$, or $\equiv 2\pmod{4}$ with $p\mid V$.
(3)For every prime number $p|N_{F/\mathbb{Q}}(S)$ with $p\equiv 3\pmod{4}$ and $p|abc$, say $p|a$ and $p|b$, and $(\frac{c}{p})=1$, conditions in Lemma 2.2 are satisfied.
(4)If at least one of $a,b,c$ is $\equiv 1\pmod{8}$, say, $a\equiv 1(mod8)$,
(4-1)For $b\not\equiv 1\pmod{4}$, there is an $S^{\prime}\in O_{F}$ satisfying $VS=2^{f}(2S^{\prime}+1)$, $f\in\mathbb{Z}$;
(4-2)For $b\equiv 5\pmod{8}$, $s_{0}+s_{1}A+s_{2}e(b)+s_{3}e(c)=2^{f}S^{\prime\prime}$, where $VS^{\prime\prime}\equiv 1,3,\frac{\pm 3\pm\sqrt{5}}{2}\pmod{4}$, $f=v_{2}(S)$, ;
(4-3)For $b\equiv 1\pmod{8}$, $s_{0}+s_{1}A+s_{2}e(b)+s_{3}e(c)=2^{f}S^{\prime\prime}$, where $VS^{\prime\prime}\equiv 1\pmod{4}$, $f=v_{2}(S)$.
(5)For every prime number $p\equiv 3\pmod{4}$ with $p\nmid abcN_{F/\mathbb{Q}}(S)$, $p\nmid V$.
Proof.
Directly verify that $(\frac{-1,VS}{*})$ are all 1, comparing with Theorem 2.3. ∎
Remark.
The theorem above, together with Lemma 2.1, 2,2 and their following remarks, making the process of determining a number in a biquadratic field can be finished on computer.
Example 2.5.
Let $S=7+2\sqrt{2}+2\sqrt{3}+\sqrt{6}$ in $K=\mathbb{Q}(\sqrt{2},\sqrt{3})$. $2O_{K}$ is totally ramified, hence $K$ has only one dyadic spot. One can easily verify that $S$ is totally positive, and $N_{K/\mathbb{Q}}(S)=1009$, a prime $\equiv 1\pmod{4}$. Hence $S$ is a sum of two squares at all spots, and therefore globally. Actually,
$S=(1-\frac{\sqrt{2}}{2})^{2}+(1+\frac{\sqrt{6}}{2}+\sqrt{3})^{2}.$
Example 2.6.
Let $S=2+\sqrt{-3}+\sqrt{5}-3\sqrt{-15}$ in $K=\mathbb{Q}(\sqrt{-3},\sqrt{5})$. $O_{K}$ has two dyadic spot isomorphic to $\mathbb{Q}_{2}(\sqrt{5})$. $N_{K/\mathbb{Q}}(S)=20629=7^{2}\times 421$, where $421\equiv 1\pmod{4}$ and $7\equiv 3\pmod{4}$. Since $(\frac{3}{7})=1$, $(\frac{5}{7})=-1$, 7-adic primes are isomorphic to $\mathbb{Q}_{7}(\sqrt{-1})$, so no verification is needed in 7-adic cases. As for dyadic cases,
$S^{\prime\prime}=2\sqrt{5}-1=4(\frac{\sqrt{5}-1}{2})+1\equiv 1\pmod{4}.$
Hence $S$ is the sum of two squares at both dyadic places, therefore globally. Actually,
$S=(\frac{5}{2}-\sqrt{5}+\frac{\sqrt{-15}}{2}+2\sqrt{-3})^{2}+(3\sqrt{-3}-\frac{3}{2}+\frac{\sqrt{-15}}{2}-2\sqrt{5})^{2}.$
Example 2.7.
Let $S=-92+\sqrt{-7}+21\sqrt{17}-\sqrt{-119}$ in $K=\mathbb{Q}(\sqrt{-7},\sqrt{17})$. $K$ has 4 dyadic spots, all of which isomorphic to $\mathbb{Q}_{2}$. $N_{K/\mathbb{Q}}(S)=3130541$, a prime $\equiv 1\pmod{4}$. So we only need to focus on dyadic cases. We have $e(-7)=53$ and $e(17)=23$, so $S^{\prime\prime}=-92+53+21\times 23-23\times 53=-775\equiv 1\pmod{8}$. Hence $S$ is a sum of two squares in $K$. Actually,
$S=(\frac{3}{2}+\sqrt{-7}-\frac{\sqrt{-119}}{2})^{2}+(\frac{1}{2}-2\sqrt{-7}+\frac{\sqrt{-119}}{2})^{2}.$
Acknowledgements
The author is sincerely grateful for directions by Prof Hourong Qin, and support by National Natural Science Foundation of China (11971224).